# Rogers’s proof of Vaaler’s theorem

### v2

@@ Acknowledgments
The author thanks Grigory M. Ivanov for the discussions and suggestions, and the anonymous referee for pointing to some language issues.
@@ 2. Proof of the volume estimate
(the index denotes the codimension) consider the simplex $a_{0}a_{1}\dots a_{n}$, settings $a_{k}=a_{F_{k}}$ for brevity and observe (by induction on the skeleta of $F$) that these simplices do not overlap and cover $P$. The situation is similar to a barycentric subdivision, but here a point $a_{F}$ need not lie in the relative interior of $F$, hence some simplices of the subdivision may be degenerate.
Now we consider one such non-degenerate simplex $A=\operatorname{conv}\{a_{0},a_{1},\ldots,a_{n}\}$ and also consider another simplex $B=\operatorname{conv}\{b_{0},b_{1},\ldots,b_{n}\}$, where $b_{k}$ is the closest to the origin point in the affine span $\operatorname{aff}F_{k}$. By the assumption of the theorem $|b_{k}|\geq\sqrt{k}$ for any $k$. The simplex $B$ has the property that its edges $b_{k}b_{k+1}$ are pairwise orthogonal, which is easy to see in the orthogonal coordinate system in which the flag of affine subspaces $\operatorname{aff}F_{k}$ is the flag of coordinate subspaces.
@@
$\frac{\operatorname{vol}B_{0}(r)\cap A}{\operatorname{vol}A}=\frac{f(\operatorname{vol}B_{0}(r)\cap A)}{\operatorname{vol}C}\leq\frac{\operatorname{vol}B_{0}(r)\cap C}{\operatorname{vol}C}.$
@@
Let us use the terminology for simplices like $B$ in the above proof.
@@
Set $\lambda_{k}=\gamma_{k}/\beta_{k}>0$. The map $f:B\to C$ in these coordinates is the linear transformation by a diagonal matrix with $\lambda_{1},\ldots,\lambda_{n}>0$ on the diagonal. Let us show that under this diagonal transformation $|x|$ cannot increase for any $x\in B$. Assume that
$x=(\beta_{1}x_{1},\ldots,\beta_{m}x_{n})\in B\Rightarrow 1\geq x_{1}\geq x_{2}\geq\dots\geq x_{n}\geq 0,$
then its image is
$y=(\gamma_{1}x_{1},\ldots,\gamma_{m}x_{n}).$
The inequality $|y|^{2}\leq|x|^{2}$ in coordinates reads
@@
This is a linear inequality on the vector $(x_{1}^{2},x_{2}^{2},\ldots,x_{n}^{2})$ that satisfies the system of linear inequalities roughly defining $B$:
@@ 3. Proof of the surface area estimate
If one wants to bound the surface area of $P$ (the (n1)(n-1)-dimensional Riemannian volume of $\partial P$) from below then one may observe that $A$ and $B$ have the same distance from the origin to the affine spaces of opposite facets, hence

### v3

@@ Abstract.
The research was carried out within the state assignment 1.1.1-0029/25 of Ministry of Science and Higher Education of the Russian Federation for IITP RAS
@@ Acknowledgments
The author thanks Grigory M. Ivanov for the discussions and suggestions, and the anonymous referees for carefully reading the text and suggesting several corrections.
@@ 2. Proof of the volume estimate
(the index denotes the codimension) consider the simplex $a_{0}a_{1}\dots a_{n}$, setting $a_{k}=a_{F_{k}}$ for brevity and observe (by induction on the skeleta of $F$) that these simplices do not overlap and cover $P$. The situation is similar to a barycentric subdivision, but here a point $a_{F}$ need not lie in the relative interior of $F$, hence some simplices of the subdivision may be degenerate.
Now we consider one such non-degenerate simplex $A=\operatorname{conv}\{a_{0},a_{1},\ldots,a_{n}\}$ and also consider another simplex $B=\operatorname{conv}\{b_{0},b_{1},\ldots,b_{n}\}$, where $b_{k}$ is the point of the affine span $\operatorname{aff}F_{k}$ closest to the origin. By the assumption of the theorem $|b_{k}|\geq\sqrt{k}$ for any $k$. The simplex $B$ has the property that its edges $b_{k}b_{k+1}$ are pairwise orthogonal, which is easy to see in the orthogonal coordinate system in which the flag of affine subspaces $\operatorname{aff}F_{k}$ is the flag of coordinate subspaces.
@@
$\frac{\operatorname{vol}B_{0}(r)\cap A}{\operatorname{vol}A}=\frac{\operatorname{vol}f(B_{0}(r)\cap A)}{\operatorname{vol}C}\leq\frac{\operatorname{vol}B_{0}(r)\cap C}{\operatorname{vol}C}.$
@@
Let us use the standard terminology for simplices like $B$ in the above proof.
@@
Set $\lambda_{k}=\gamma_{k}/\beta_{k}>0$. The map $f:B\to C$ in these coordinates is the linear transformation by a diagonal matrix with $\lambda_{1},\ldots,\lambda_{n}>0$ on the diagonal. Let us show that under this diagonal transformation $|x|$ cannot increase for any $x\in B$. A point
$x=(\beta_{1}x_{1},\ldots,\beta_{m}x_{n})$
is in $B$ if and only if
$x_{1}\geq x_{2}\geq\dots\geq x_{n}\geq 0.$
Its image is
$f(x)=(\gamma_{1}x_{1},\ldots,\gamma_{m}x_{n})$
and the inequality $|f(x)|^{2}\leq|x|^{2}$ in coordinates reads
@@
This is a linear inequality on the vector $(x_{1}^{2},x_{2}^{2},\ldots,x_{n}^{2})$ that satisfies the system of linear inequalities following from those defining $B$:
@@ 3. Proof of the surface area estimate
If one wants to bound the surface area of $P$ (the (n1)(n-1)-dimensional Riemannian volume of $\partial P$) from below then one may observe that $A$ and $B$ have the same distance from the origin to the affine spans of opposite to the origin facets, hence